# AI Agentic Programming: A Survey of Techniques, Challenges, and Opportunities

### v1

@@ 6.6. Safety, Alignment, and Trust
As agentic coding systems become more autonomous and start taking on greater responsibility in software development, it becomes increasingly important to ensure that their behavior aligns with user intent, produces correct results, and avoids unintended changes (ouyang2022training, ; liu2024exploring, ). Unlike traditional LLM-based code assistants, agentic systems can take multi-step actions, use external tools, and modify codebases with limited human oversight, making safety and trust essential goals (nvidia2025safetyrecipe, ; hua2024trustagent, ).
One key research direction is building agents that can better understand and follow what users actually want, even when instructions are vague or incomplete. Current systems rely heavily on natural language prompts, which can be ambiguous. Future work could focus on grounding agent behavior in more structured forms of input, such as constraints, test cases, or high-level goals, that are easier to validate and reason about (fakhoury2024llm, ; huang2023agentcoder, ; li2025structured, ).
@@
An interesting idea is to design a structured language that developers can use to express their intent clearly. This language would serve as a kind of programming interface for LLMs, allowing users to define what the agent should do, what it must avoid, and what counts as a valid solution. For example, a developer might specify that a function’s output must remain the same after refactoring, and the agent would only explore changes that meet this requirement. This kind of structure could also make it easier to apply lightweight verification tools, such as static analyzers or type checkers, to ensure the agent’s suggestions are safe and correct (codeql2020, ; li2024iris, ; nunez2024autosafecoder, ).
In safety-critical domains like healthcare, finance, or automotive software, agents will also need to follow strict coding standards, legal rules, and certification guidelines. Agents could be trained to recognize such constraints or be paired with rule-based systems to flag violations and suggest compliant alternatives (ouyang2022training, ; he2024instruction, ).
@@
Another important challenge is ethical alignment. Because agentic systems are trained on large and diverse codebases, they may learn unsafe or biased practices (liu2023uncovering, ; pearce2025asleep, ; huang2024bias, ). Research efforts are needed to detect and reduce risks such as generating insecure code, leaking sensitive data, or reinforcing stereotypes. Techniques from responsible AI, like behavioral audits, adversarial testing, and human feedback, can be adapted to this setting.
To build user trust, agents should also be able to explain their decisions (rajani2019explain, ; creswell2022faithful, ; DBLP:conf/nips/ShinnCGNY23, ). Developers need to know why an agent made a change, where the idea came from, and what trade-offs are involved. Research on explanation generation, source citation, and visualization of code changes will help make agent behavior more transparent and understandable (liu2024reliability, ; sun2022investigating, ; huang2023citation, ).
Agents should also be aware of their own limitations. When they are uncertain, they should be able to flag their confidence level (kuhn2023semantic, ; creswell2022faithful, ), suggest multiple options (DBLP:conf/nips/MadaanTGHGW0DPY23, ; DBLP:conf/nips/ShinnCGNY23, ), or ask the user for confirmation (mu2024clarifygpt, ; zhang2024modeling, ). Designing agents that can adapt their behavior based on task difficulty or user feedback is a promising direction (DBLP:conf/nips/ShinnCGNY23, ; sun2023adaplanner, ; damani2024learning, ; manvi2024adaptive, ; chen2024magicore, ).
Finally, safety mechanisms like undo and audit trails will be essential. Agents should keep track of what they have changed and allow users to roll back actions easily (patil2024goex, ; jiang2024rocode, ). Future work could explore automatic snapshotting, reversible code edits, and tight integration with version control systems to support safe collaboration.
@@ 7. Conclusion
We have presented a comprehensive review of AI agentic programming. This new software paradigm, driven by the success of large language models (LLMs), is a transformative shift in how software can be created, maintained, and evolved. By combining the capabilities of LLMs with planning, tool use, and iterative refinement, coding agents are beginning to automate complex, multi-stage programming workflows that traditionally required significant human involvement. These systems can not only generate and test code but also interact with development tools, decompose tasks, and adapt based on feedback, bringing us closer to the vision of autonomous software development.
In this article, we have introduced a taxonomy of AI coding agents and architectures, reviewed underpinning techniques like context management and tool integration, and discussed how existing benchmarks evaluate the capabilities of coding agents. We have summarized the progress made in enabling LLM-based AI agents to reason over tasks, interface with software development tools, and operate in increasingly sophisticated ways. At the same time, we have identified multiple open challenges that must be addressed to ensure these systems are safe, reliable, and usable in real-world settings. These include limitations in context handling, the need for persistent and structured memory, alignment with user intent, human-AI collaboration, and verification of agent behavior. As software developers are increasingly relying on AI coding agents, these concerns will become more pressing and will require interdisciplinary solutions drawing from programming languages, human-computer interaction, software engineering, and responsible AI.

### v2

@@ 6.6. Safety, Alignment, and Trust
As agentic coding systems start taking on greater responsibility in software development, it becomes increasingly important to ensure that their behavior aligns with user intent, produces correct results, and avoids unintended changes (Ouyang et al., 2022; Liu et al., 2024b). Unlike traditional LLM-based code assistants, agentic systems can take multi-step actions, use external tools, and modify codebases with limited human oversight, making safety and trust essential goals (NVIDIA Technical Blog, 2025; Hua et al., 2024).
@@
One key research direction is building agents that can better understand and follow what users actually want, even when instructions are vague or incomplete. Current systems rely heavily on natural language prompts, which can be ambiguous. Future work could focus on grounding agent behavior in more structured forms of input, such as constraints, test cases, or high-level goals, that are easier to validate and reason about (Fakhoury et al., 2024; Huang et al., 2023; Li et al., 2025b). Another interesting idea is to design a structured language that developers can use to express their intent clearly. This language would serve as a kind of programming interface for LLMs, allowing users to define what the agent should do, what it must avoid, and what counts as a valid solution. For example, a developer might specify that a function’s output must remain the same after refactoring, and the agent would only explore changes that meet this requirement. This kind of structure could also make it easier to apply lightweight verification tools, such as static analyzers or type checkers, to ensure the agent’s suggestions are safe and correct (Lab, 2020; Li et al., 2024a; Nunez et al., 2024).
@@
In safety-critical domains like healthcare, finance, or automotive software, agents will also need to follow strict coding standards, legal rules, and certification guidelines. Agents could be trained to recognize such constraints or be paired with rule-based systems to flag violations and suggest compliant alternatives (Ouyang et al., 2022; He et al., 2024a). Another important challenge is ethical alignment. Because agentic systems are trained on large and diverse codebases, they may learn unsafe or biased practices (Liu et al., 2023a; Pearce et al., 2025; Huang et al., 2024g). Research efforts are needed to detect and reduce risks such as generating insecure code, leaking sensitive data, or reinforcing stereotypes. Techniques from responsible AI, like behavioral audits, adversarial testing, and human feedback, can be adapted to this setting.
To build user trust, agentic programming systems should also be able to explain their decisions (Rajani et al., 2019; Creswell and Shanahan, 2022; Shinn et al., 2023). Developers want to know why an agent made a change, where the idea came from, and what trade-offs are involved. Research on explanation generation, source citation, and visualization of code changes will help make LLM behavior more transparent and understandable (Liu et al., 2024d; Sun et al., 2022; Huang and Chang, 2023). Agents should also be aware of their limitations. When they are uncertain, they should be able to flag their confidence level (Kuhn et al., 2023; Creswell and Shanahan, 2022), suggest multiple options (Madaan et al., 2023; Shinn et al., 2023), or ask the user for confirmation (Mu et al., 2024; Zhang et al., 2024c). Designing agents that can adapt their behavior based on task difficulty or user feedback is a promising direction (Shinn et al., 2023; Sun et al., 2023b; Damani et al., 2024; Manvi et al., 2024; Chen et al., 2024d). Furthermore, safety mechanisms like undo and audit trails will be essential. Agents should keep track of what they have changed and allow users to roll back actions easily (Patil et al., 2024; Jiang et al., 2024a). Future work could explore automatic snapshotting, reversible code edits, and tight integration with version control systems to support safe collaboration.
6.7. Multi-agent Collaboration
As coding agents increasingly operate in teams and interact with heterogeneous services, effective multi-agent collaboration requires standardized communication protocols that provide low-latency, scalable, and semantics-aware coordination. However, current solutions remain limited and fall short of these needs. Future communication protocols should move beyond ad-hoc adapters and heterogeneous web services toward native support for multi-agent collaboration. Promising directions include the design of lightweight, low-latency group communication frameworks that natively support both agent–to–agent and agent–to–tool interactions, while offering standardized session management for dynamic participation and coordination (Xu et al., 2012; A2A, 2025; Verma et al., 2015). Recent standards efforts such as IEEE P3394 propose a universal framework with session, channel, and message abstractions, enabling consistent semantics and interoperability across diverse transports (Tong et al., 2025). Building on these initiatives, future research should emphasize semantics-aware routing, conflict resolution, and consistency guarantees in distributed multi-agent settings. Ultimately, advancing collaboration-centric protocols will be essential for scaling efficient agentic systems in practice.
6.8. System Support for AI Coding Agents
Realizing agentic programming at scale also requires strong systems-level support. When offered as a cloud service (Amazon, 2025; Microsoft, 2025; KServe Community, 2025; Hugging Face, 2025), LLMs and coding agents must be backed by infrastructures that can efficiently manage resources, communication, and state. Agentic workflows combine code generation, compilation, testing, and tool orchestration, with highly variable resource demands (Yang et al., 2024b; Wang et al., 2024d). Systems must schedule CPU, GPU, and memory adaptively, balance throughput across users, and minimize latency for interactive tasks (Yu et al., 2021; Fu et al., 2024; Yu et al., 2022; Patel et al., 2024; Aminabadi et al., 2022). During the iterative code generation and testing process, agents are likely to exchange large volumes of data with tools like compilers, debuggers, profilers, and repositories. Reducing overhead through caching (Kwon et al., 2023; Liu et al., 2024a), compression (Yang et al., 2024a; Jiang et al., 2023; Hooper et al., 2024; Zhao et al., 2024), and lightweight protocols (Google, 2015; Moritz et al., 2018) is essential for efficiency in distributed environments. Equally important is context and state management. To support interruption, resumption, and collaboration, systems must persist not only source code but also goals, annotations, and intermediate representations. Efficient checkpointing and synchronization are key enablers here. Finally, multi-agent collaboration introduces distributed systems challenges: coordinating tasks, resolving conflicts, and ensuring consistency when multiple agents - or human–agent teams - work concurrently on shared codebases.
@@ 7. Conclusion
We have presented a comprehensive review of AI agentic programming. This new software paradigm, driven by the recent success of LLMs, is a transformative shift in how software can be created, maintained, and evolved. By combining the capabilities of LLMs with planning, tool use, and iterative refinement, coding agents are beginning to automate complex, multi-stage programming workflows that traditionally required significant human involvement. These systems can not only generate and test code but also interact with development tools, decompose tasks, and adapt based on feedback, bringing us closer to the vision of autonomous software development.
We have introduced a taxonomy of AI coding agents and architectures, reviewed underpinning techniques like context management and tool integration, and discussed how existing benchmarks evaluate the capabilities of coding agents. We have summarized the progress made in enabling LLM-based AI agents to reason over tasks, interface with software development tools, and operate in increasingly sophisticated ways. At the same time, we have identified multiple open challenges that must be addressed to ensure these systems are safe, reliable, and usable in real-world settings. These include limitations in context handling, the need for persistent and structured memory, alignment with user intent, human-AI collaboration, and verification of agent behavior. As software developers are increasingly relying on AI coding agents, these concerns will become more pressing and will require interdisciplinary solutions drawing from programming languages, human-computer interaction, software engineering, and responsible AI.